### v3

Abstract
Many studies have shown how ergonomically designed furniture improves productivity and well-being. As computers have become a part of students’ academic lives, they will grow further in the future. We propose anthropometric-based furniture dimensions suitable for university students to improve computer laboratory ergonomics. We collected data from 380 participants and analyzed 11 anthropometric measurements, correlating them to 11 furniture dimensions. Two types of furniture were found and studied in different university computer laboratories: (1) a non-adjustable chair with a non-adjustable table and (2) an adjustable chair with a non-adjustable table. The mismatch calculation showed a significant difference between existing furniture dimensions and anthropometric measurements, indicating that 7 of 11 existing furniture dimensions need improvements. The one-way ANOVA test with a significance level of 5% also showed a significant difference between anthropometric data and existing furniture dimensions. All 11 dimensions were determined to match the student’s anthropometric data. The proposed dimensions were found to be more compatible and showed reduced mismatch percentages for 9 furniture dimensions and nearly zero mismatches for SW, BH, and UTH for both males and females compared to existing furniture dimensions. The proposed dimensions of the furniture set with adjustable seat height showed slightly improved results for SH and STC, which showed zero mismatches compared to the non-adjustable furniture set. The TW and TD dimensions were suggested according to Barnes and Squires’ ergonomic work envelope model, considering hand reach. The position for the keyboard and mouse was also suggested according to the work envelope. The monitor position and viewing angle were proposed according to OSHA guidelines. This study suggests that the proposed dimensions can improve comfort levels, reducing the risk of musculoskeletal disorders among students. Further studies on the implementation and long-term effects of these proposed dimensions in real-world computer laboratory settings are recommended.
keywords:
Ergonomics, ANOVA, Anthropometric measurements, Mismatch analysis, Computer lab furniture, Furniture design
[…]
[inst1]organization=Department of Industrial Engineering and Management,addressline=Khulna University of Engineering and Technology (KUET), city=Khulna, postcode=9203, country=Bangladesh
[…]
[inst2]organization=Department of Computer Science,addressline=American International University-Bangladesh (AIUB), city=Dhaka, postcode=1229, country=Bangladesh
1 Introduction
The use of computers in professional environments is experiencing an upward trend. Moreover, the use of home computers for both occupational and educational purposes has seen a notable surge over the preceding decades (Py Szeto et al.,, 2014). Poor posture linked to furniture is a recognized factor that can increase the likelihood of developing musculoskeletal disorders (MSDs) (Magnusson and Pope,, 1998). Therefore, human body measurements are the most critical factor in designing furniture that can effectively adapt to posture, thereby enhancing comfort (Kopardekar,, 1996). Anthropometry plays a vital role in designing work environments to reduce the occurrence of MSDs among employers and employees, ultimately leading to increased productivity (Afzan et al.,, 2012). A study was conducted on the postural angles of primary body segments, including the neck, elbow, and wrist, among six sets of parents and children while using a computer at home (Williams and Jacobs,, 2002). Ergonomics seeks to create work environments that allow individuals to work easily, efficiently, and safely (Savanur et al.,, 2007). Creating equipment, merchandise, furniture, and other devices that match the human body’s needs can enhance the comfort, physical health, well-being, and performance of individuals (Parvez et al.,, 2018).
The motivation for this study stems from the growing recognition of the importance of ergonomic design in enhancing human productivity and well-being. In recent years, numerous studies have demonstrated the positive impact of ergonomically designed furniture on comfort, health, and performance, particularly in settings where individuals spend prolonged periods of time seated at desks, such as offices and educational institutions. As computers and digital technology continue to permeate various aspects of academic and professional life, ensuring optimal ergonomic conditions in computer labs has become increasingly crucial. While ergonomic guidelines and standards exist for general office environments, there is a notable gap in tailored guidelines for computer labs, particularly for university students.
The prevalence of computers in educational settings, especially universities, highlights the need to prioritize students’ physical comfort and health by providing ergonomically sound furniture. However, the lack of specific guidelines for computer laboratories presents a challenge in achieving this objective. This study addresses this gap by proposing anthropometric-based furniture dimensions that are specifically designed for university students. By analyzing data collected from 380 participants and correlating 11 anthropometric measurements to 11 furniture dimensions, this research aimed to optimize furniture design in university computer labs to better accommodate students’ physical characteristics.
We aimed to identify potential mismatches between students’ anthropometric measurements and computer lab furniture dimensions to establish better dimensions for the design parameters of computer lab furniture for university students, considering relevant anthropometric measurements of students sitting to minimize potential mismatches and prolonged health issues. This research aims to assess and compare the potential mismatch between two existing sets of computer tables and chairs and propose standardized design criteria and dimensions based on anthropometry and safety guidelines for students. This will involve analyzing 11 anthropometric data points along 11 furniture dimensions using ANOVA for decision-making. It is anticipated that aligning furniture more closely with the human body will result in a notable reduction in MSDs and an associated increase in productivity. Both non-adjustable and adjustable furniture dimensions must be adjusted according to the anthropometric data of the undergraduate students.
This study focuses on optimizing university furniture design using personalized anthropometric measurements tailored to university student populations. By analyzing 11 key anthropometric measurements and comparing them with existing furniture, I provide a comprehensive understanding of students’ ergonomic needs. ANOVA statistical analysis adds robust validation to my findings, ensuring that any proposed ergonomic design is grounded in solid statistical evidence. My research proposes innovative and ergonomically designed furniture solutions based on current student body measurement data, potentially improving student comfort, productivity, and overall well-being. These findings have significant implications for educational institutions and furniture manufacturers, highlighting the importance of customizing furniture designs to meet users’ specific needs. My research contributes to the ergonomics and design fields by addressing the literature gap regarding the ergonomic needs of university students. The proposed ergonomic designs have practical applications in real-world settings, such as university classrooms, computer labs, and study areas, where they can enhance the learning environment and promote student health and well-being. This research also opens the door for further exploration of the ergonomic needs of different user groups and the development of more tailored furniture solutions for various settings.
The contributions of the proposed work are given below:
  1. 1.
    This study highlights the importance of considering anthropometric measurements in furniture design for educational institutions, serving as a reference for future research and informing the design of ergonomic furniture in other educational settings.
  2. 2.
    By identifying the mismatch between existing furniture dimensions and anthropometric measurements, this research offers practical insights for designing ergonomic furniture in university environments, improving comfort, health, and productivity among students and staff.
  3. 3.
    Designing ergonomic furniture that meets the specific needs of university students and staff may lead to potential cost savings for educational institutions through reduced absenteeism, increased productivity, and improved overall well-being.
  4. 4.
    We collected and analyzed anthropometric data from university students, providing valuable insights into the body measurements of the target population.
  5. 5.
    We conducted a correlation analysis among 11 anthropometric measurements, shedding light on the interrelationships between different body dimensions.
  6. 6.
    By comparing the mismatch between two different types of existing furniture and anthropometric data, the study offers a nuanced understanding of how different furniture designs affect users’ comfort and well-being.
  7. 7.
    The study employed ANOVA for better decision-making, providing a robust statistical framework for assessing the significance of differences among existing furniture types and anthropometric measurements.
  8. 8.
    Drawing from the current students’ body measurement data, this study proposes ergonomically designed furniture that aligns with the specific needs and dimensions of the target population, offering practical solutions for enhancing comfort and productivity in university settings.
The paper is structured as follows: The “Introduction” section overviews the research problem and outlines the significance of addressing ergonomic challenges in university environments. In the “Literature Review” section, we review existing studies on ergonomics and anthropometric measurements in university settings, highlighting the gaps that our research aims to fill. The “Methodology” section details our approach to collecting anthropometric data and designing furniture based on these measurements. In the “Results and Discussion” section, we present the findings of our anthropometric survey, evaluate the proposed ergonomic furniture design, interpret and discuss the results. The “Conclusions” section summarizes the main findings and contributions of the study and proposes future research directions.
2 Literature Review
Self et al., (2019) conducted a study that examined the potential factors that promote or hinder interdisciplinary education in design, specifically from the viewpoint of student learning. The researchers conducted a series of detailed interviews with students from two interdisciplinary design schools, one located in the Republic of Korea and the other in the UK. This study discusses the implications of adopting interdisciplinary approaches in design education and proposes areas for further research in this domain (Self et al.,, 2019). Laeser et al., (1998) recognized the necessity for furniture designed for 6th and 8th grade children. In their study, students from these grades performed keyboarding and mousing tasks at two distinct computer workstations. One workstation was a conventional desktop setup, whereas the other featured an adjustable keyboard. The study found that the overall posture scores of the students improved when using the adjustable workstation, as indicated by the Rapid Upper Limb Assessment. Thariq, (2010) found that current side-mounted desktop chair designs are not suited to fulfilling the postural and comfort requirements of university students in their learning environment. In response, the researchers conducted an anthropometric survey of university students in Sri Lanka. The survey aimed to create fixed-type and side-mounted desktop chairs and evaluate their comfort levels based on design dimensions. This research emphasizes the significance of chairs that support the development of a healthy and comfortable posture among students, which can contribute to creating a healthy workforce in various industries. Taifa and Desai, (2017) recognized the issue of poorly designed furniture for engineering students in India. The authors employed a health survey, specifically an ergonomic assessment, to gather and analyze the data in their study. Based on their findings, the authors proposed the design of adjustable classroom furniture as a solution to reduce MSDs.
Hoque et al., (2014) found that classroom furniture does not consider ergonomics. The authors conducted an anthropometric survey of 500 Bangladeshi university students, followed by a mismatch analysis. Based on their findings, the authors recommend the use of ergonomically fit furniture. Joshi et al., (2015) addressed the significant rise in the design of computer workstations and associated health and safety issues for users. Their research aimed to evaluate students’ knowledge at the G.B.P. University of Agriculture and Technology concerning the use of ergonomically appropriate computer workstations. The authors conducted a postural assessment using the participative ergonomic technique of Rapid Upper Limb Assessment (RULA). The study revealed that most computer users lacked adequate information regarding the positioning of computer workstations, and a substantial number of computer operators reported issues with their upper extremities. Mowatt et al., (2018) aimed to determine the prevalence of computer vision syndrome (CVS) and ergonomic practices among students in the Faculty of Medical Sciences at The University of the West Indies, Jamaica. A cross-sectional study was conducted for research purposes using a self-administered questionnaire. The study found that ocular symptoms and neck pain were less likely when the device was held just below eye level. The study highlighted a high prevalence of CVS symptoms among university students, suggesting that adopting improved ergonomic practices could potentially reduce issues such as neck pain, eye strain, and burning. Kahya, (2019) investigated the congruence between the dimensions of school furniture and anthropometric measurements of students. They assessed nine anthropometric dimensions of 225 students from nine departments. The results indicated a significant mismatch, with a seat height discrepancy of 44.45% discrepancy, seat depth of 100% discrepancy, and desk height of 21.28% discrepancy. This study lays the groundwork for and encourages further exploration of classroom ergonomics in university settings. Abd Rahman et al., (2018) developed an anthropometric database tailored to Malaysian operators, focusing on sitting and standing dimensions. The database was constructed from measurements of 146 male and 168 female participants aged 18 - 45 years. Thirty-six anthropometric dimensions were chosen to compare their common availability across the four countries. This knowledge of population variations is crucial for designing workstations and facilities that accommodate the needs of the industrial environment.
Zayed et al., (2021) sought to understand the prevalence and factors associated with Digital Eye Strain (DES) among IT professionals at Tanta University, Egypt. They employed an interview questionnaire to collect and analyze data, revealing significant predictors of DES, including sex, age, wearing eyeglasses, prolonged computer use, an unadjusted ergonomic workstation, and a dry environment. The study concluded that DES can be mitigated by increasing knowledge and awareness through eye health education, regular training in ergonomic workstations, and adjustments to create a comfortable workplace environment. Parvez et al., (2018) identified that poorly fitting furniture could contribute to various MSDs and discomforts. To delve into this issue, they surveyed ten primary schools in Bangladesh, revealing a notable disparity between furniture dimensions and student body measurements. Moreover, this paper proposes furniture dimensions that could notably decrease the mismatch percentage among students from 90% to 10%. Shohel Parvez et al., (2022) acknowledged that inadequately designed furniture could potentially exacerbate MSDs. They conducted a survey of 400 university students (250 males and 150 females) who volunteered for the study. The Standard Nordic Musculoskeletal Questionnaire was used to evaluate MSDs. The authors recommended that modifying the design of academic furniture could help alleviate or prevent MSD.
Dessie et al., (2018) observed that prolonged computer use could increase the risk of CVS in users. They conducted a cross-sectional study to gauge the prevalence of CVS and its related factors among government employees in Debre Tabor town, encompassing 607 participants. The research revealed that computer users with a comprehensive understanding of computer use safety measures and associated adverse effects were less susceptible to CVS. Boadi-Kusi et al., (2020) emphasized the importance of addressing poor ergonomic office setups that can lead to university administrative staff in Ghana spending prolonged periods behind computer screens, thus increasing the risk of developing CVS. To address this issue, a cross-sectional survey was conducted among 200 administrative staff members at the University of Cape Coast. The results indicated that raising awareness about CVS and adhering to recommended ergonomic practices are crucial in reducing the prevalence of CVS and, consequently, enhancing work satisfaction and productivity. Moradi et al., (2017) conducted an ergonomic risk assessment of auto mechanics. This study found that the back and waist had the highest prevalence of work-related MSDs. Fazi et al., (2019) utilized the Rapid Upper Limb Assessment (RULA) method to assess risks in an automotive manufacturing company. These findings demonstrated an exceptionally high level of risk, highlighting the urgent need for immediate changes. Kibria, (2019) developed a self-report Nordic ergonomic assessment questionnaire to identify discrepancies between furniture dimensions and anthropometric measurements. The proposed ergonomically designed computer workstation, informed by anthropometric measurements and guidelines, could potentially alleviate MSDs among university teachers. This study could significantly influence ergonomic furniture design in universities and other organizations, addressing computer-related ergonomic challenges.
The current literature on ergonomics and its impact on user health and performance has identified several research gaps that warrant further investigation. First, comprehensive research on interdisciplinary education in design and its impact on student learning outcomes is lacking. Second, few studies have been conducted on the impact of ergonomic furniture design on children’s posture and health. Third, there is a shortage of research on the ergonomic needs of university students, particularly in different countries and cultural contexts. Fourth, studies on the prevalence and impact of CVS among university students and professionals in various occupational settings are scarce. Fifth, there is a lack of research on the effectiveness of ergonomic interventions in reducing the risk of MSDs and other ergonomic-related issues in occupational settings. Sixth, limited research has been conducted on the impact of ergonomic interventions on improving work satisfaction, productivity, and overall well-being in various occupational settings. Seventh, there is a shortage of research on the long-term effects of ergonomic interventions on employees’ health, performance, and retention. Finally, there is a lack of studies on the cost-effectiveness of ergonomic interventions and their potential return on investment in different occupational settings. Addressing these research gaps will contribute to developing more effective ergonomic interventions and overall well-being in various occupational settings. However, there has been a limited emphasis on the design of furniture and computer labs for undergraduate education in universities, and there is a lack of established standards for computer lab furniture in Bangladeshi universities. Bangladeshi universities typically use desktops, chairs, and tables as computer lab furniture. Prolonged use of these improperly designed furniture pieces could exacerbate MSDs in young adults. Hence, ergonomically fitted furniture is crucial. Failure to address these issues in a timely manner could result in undetected severe health problems among the students.
3 Methodology
The main goal of this study is to assess the current state of computer laboratories in Bangladesh. This will be accomplished by gathering anthropometric data from the students. The objective was to establish standardized criteria and measurements for the design of computer tables in Bangladesh. The sequential steps of this study are outlined in Figure 1.
Figure 1: Flowchart illustrating the step-by-step methodology followed in this research.
3.1 Problem Identification
In our research, we sought to assess the ergonomic risks that students face due to prolonged sitting at their desks and the potential development of Work-related MSDs (WMSDs). We developed a comprehensive self-reported questionnaire based on the Nordic Musculoskeletal Questionnaire (NMQ), which has been extensively used and validated in previous studies. The questionnaire was designed to gather information on the frequency and severity of musculoskeletal symptoms experienced by teachers, focusing on areas such as the neck, shoulders, back, and wrists, which are commonly affected by prolonged sitting and poor ergonomics. Participants were asked to report the frequency of musculoskeletal symptoms they experienced, classifying them as “constantly” (most of the time during the day), “occasionally” (two to four times a month), or “frequently” (more than four times a month). We also asked the participants whether they had experienced musculoskeletal pain in the past 12 months. This allowed us to assess the prevalence of musculoskeletal pain among teachers and to identify any data patterns or trends. Approximately 78.42% of students reported experiencing MSD-related pains. In this group, 77.67% were male, and 81.25% were female. We collected demographic information, such as age, gender, and level of study, to better understand how these factors might influence the prevalence and severity of musculoskeletal symptoms. The questionnaire was administered to a sample of KUET students. The collected data were analyzed to identify any significant associations between ergonomic factors, such as prolonged sitting and poor posture, and the development of WMSDs. The reliability test for the NMQ survey yielded a Cronbach’s alpha value of 0.77, indicating an “acceptable” level of internal consistency. This value falls within the 95% confidence interval of 0.733 and 0.803, suggesting that if the survey were repeated with different samples, the actual value of Cronbach’s alpha would likely fall within this range 95% of the time.
3.2 Sample Collection
A subset of students was chosen from two departments at the Khulna University of Engineering & Technology (KUET), Khulna, Bangladesh. Equation 1 was used to establish the necessary sample size.
$n=\frac{N}{1+Ne^{2}}$ (1)
where, $n$ represents the sample size or population size, $N$ denotes the total population, and $e$ refers to the desired precision level set at 5% with a 95% confidence level.
According to KUET’s official database, there are currently 5,240 students enrolled, comprising 4,318 males and 922 females. This yielded a male-to-female ratio of approximately 4.68:1. According to Equation 1, a minimum of 372 or more data points should be collected. This study gathered anthropometric data from 300 males and 80 females, representing up to 7,000 students. All the students’ ages ranged from 17 to 26 years. The Director of Research & Extension, KUET, approved the experimental procedures, which were conducted in accordance with the principles stated in the Declaration of Helsinki. Participation in the survey was voluntary, and participants had the right to withdraw at any time without penalties. Confidentiality of participant responses was strictly maintained, and all data collected were anonymized and stored securely. The dataset, available for further exploration, can be found in “Anthropometric Data of KUET students(Jahin and Saha,, 2024).
Table 1: Demographic distribution of study population by gender and academic level
Study level Male Female
1st year student 70 20
2nd year student 70 20
3rd year student 70 20
4th year student 90 20
Total 300 80
3.3 Anthropometric Measurements
The 11 anthropometric measurements described in Table 2 were obtained for this study (see Figure 2). These dimensions were measured using definitions employed in earlier studies (Kroemer et al.,, 1997). A chair and a Harpenden Anthropometer were used to determine the seating dimensions, and sitting assistance was used to support participants in maintaining a straight posture with their feet flat on the floor. Before data collection, rigorous calibration of all equipment was conducted in accordance with the established standards to maintain the accuracy and precision of the measurements. Throughout the trial, the participants were instructed to maintain a straight posture on a flat-surfaced seat, with their knees positioned at a $90^{\circ}$ angle. Standardized posture is crucial for the consistency and comparability of anthropometric data. Participants were briefed about the study objectives, and precise postures were maintained during anthropometric measurements. All measurements were documented in millimeters (mm).
(a) Right-side view
(b) Backside view
Figure 2: Anthropometry of the participants, (a) right-side view, (b) backside view.
Table 2: Anthropometric measurements used in this study (International Organization for Standardization,, 2017)
Anthropometric dimensions Acronym Description
Popliteal Height PH
Vertical distance from the floor to the popliteal angle, located on the lower surface of the knee, while keeping the feet flat on the floor.
Sitting Elbow Height SEH
Distance from the seat surface to the underside of the elbow, with the forearm horizontal and the upper arm hanging freely downwards.
Buttock Popliteal Length BPL
Distance from the back of the seat to the back of the knee, where the backs of the lower legs meet the underside of the thighs.
Buttock Knee Length BKL
Horizontal distance from the back of the buttocks to the back of the knee.
Hip Breadth HB
The horizontal measurement taken across the hips while the participant is sitting and keeping their knees together (maximum width).
Sitting Shoulder Height SSH
Vertical distance from the seat surface to the acromion (bony shoulder point).
Sitting Elbow Height SEB
Horizontal distance across the lateral surfaces of the elbows.
Thigh Thickness TT
Vertical distance from the seat surface to the top of the thigh, where the uncompressed soft tissue of the thigh meets the abdomen.
Arm Length AL
Vertical distance from the scapula’s lowest point (inferior angle) to the participant’s sitting surface.
Elbow Fingertip Length EFL
Horizontal distance from the back of the elbow to the tip of the middle finger in a standard sitting position.
Subscapular Height SCH
Vertical distance from the scapula’s lowest point (inferior angle) to the participant’s sitting surface.
3.4 Furniture Dimensions
The dimensions used for designing the furniture are shown in Figure 3 with identities. This study evaluated two types of university furniture used in computer labs at KUET. One was a ‘table with adjustable chairs,’ and the other was a ‘table with non-adjustable chairs,’ both commonly used in computer labs at KUET. Figure 3 depicts the computer lab furniture design parameters, whereas Figure 4 shows the two furniture sets. As such, the study measured 11 design parameters for the two types of existing computer furniture, as described in Table 3.
(a) Left-side view
(b) Top view
Figure 3: Furniture’s measurement: (a) Left-side view and (b) Top view.
(a) Table with adjustable chair
(b) Table with non-adjustable chair
Figure 4: Two types of furniture set, (a) Table with adjustable chair, (b) Table with non-adjustable chair.
Table 3: Parameters used for furniture design in this study
Furniture dimension Acronym Description
Seat Height SH Vertical distance from the floor to the top of the seat surface.
Seat Width SW Horizontal distance between the lateral edges of the seat.
Seat Depth SD Horizontal distance from the front edge to the back edge of the seat surface.
Backrest Height BH Vertical distance from the seat surface’s top edge to the backrest’s top edge.
Backrest Width BW Horizontal distance between the lateral edges of the backrest.
Seat to Table Height STH Vertical distance from the seat surface’s top edge to the table surface’s top edge.
Table Depth TD Horizontal distance from the front edge to the back edge of the table surface.
Table Length TL The maximum horizontal distance between the lateral edges of the table surface
Table Height TH Vertical distance from the floor to the top edge of the table surface.
Underneath Table Height UTH Vertical distance from the lowest point below the table to the floor’s surface.
Upper Edge of Backrest UEB Vertical distance from the backrest’s top edge to the headrest’s top edge.
3.5 Furniture Dimensions and Anthropometric Data Mismatch
Considering the students’ anthropometric measurements when designing and evaluating computer lab furniture is essential. This involves using relevant anthropometric estimations and ergonomic standards to determine the acceptable range for each furniture measurement.
3.5.1 PH against SH
Popliteal height (PH) and seat height (SH) are typically connected. SH should provide knee flexion to align the lower legs with the body’s craniocaudal axis at a $30^{\circ}$ angle (JFM et al.,, 2003). According to Parcells et al., (1999), PH should be higher than SH. This means that the shin-thigh position should be between $95^{\circ}$ and $120^{\circ}$, and the lower leg should create an angle with a vertical axis of $5^{\circ}$ to $30^{\circ}$. While an unduly low SH can strain the ischial tuberosities, an extremely high SH can increase the pressure on the bottom of the knee and nerves, decreasing blood circulation (Hoque et al.,, 2014). The SH should be less than 40 mm, or 88% of the PH, to prevent putting excessive pressure on the buttocks (JFM et al.,, 2003). The PH measurement considered the shoe height by adding 30 mm. Therefore, Equation 2 can be used to represent the relationship between PH and SH:
$(PH+30)\cos(30^{\circ})\leq SH\leq(PH+30)\cos(5^{\circ})$ (2)
3.5.2 HB against SW
To allow for both furniture flexibility and essential motions, the seat width (SW) should be broad enough to accommodate people with the widest hip breadth (Evans et al.,, 1988; Sanders and McCormick,, 1993). A proper fit is suggested by research, and the 95th percentile of HB is a good place to start when measuring SW (Evans et al.,, 1988; Occhipinti et al.,, 1993; Oborne,, 1995; Sanders and McCormick,, 1993). The SW should be wider than the HB. Previous research has suggested that the optimal SW should to be within the range of 10% to 30% of HB (Dianat et al.,, 2013; Gouvali and Boudolos,, 2006). Consequently, the link between the HB and SW can be established using Equation 3:
$(1.10\times HB)\leq SW\leq(1.30\times HB)$ (3)
3.5.3 BPL against SD
There is a strong relationship between SD and BPL, with BPL frequently serving as a guide for SD development and evaluation. The fifth percentile of the BPL should be used as the basis for the seat depth, according to several studies (Hossain and Ahmed,, 2010; Sanders and McCormick,, 1993; Oborne,, 1995; Panagiotopoulou et al.,, 2004). This guarantees that the seat backrest can support the lumbar spine without compressing the popliteal surface. The seat may not support the thigh when the SD is much shorter than the BPL (Castellucci et al.,, 2010). On the other hand, the user might be unable to use the backrest to support their lumbar spine if the SD is larger than the BPL (Milanese and Grimmer,, 2004). Accordingly, when SD is either 95% of BPL or less, there is a mismatch between the SD and BPL (Afzan et al.,, 2012; Altaboli et al.,, 2015; Dianat et al.,, 2013). Equation 4 can be used to express the connection between BPL and SD.
$(0.80\times BPL)\leq SD\leq(0.95\times BPL)$ (4)
3.5.4 SEH against STH
According to several studies, the SEH is a critical factor in SDH determination (Bendix and Bloch,, 1986; Milanese and Grimmer,, 2004; JFM et al.,, 2003). Developing an appropriate TH requires careful consideration of shoulder biomechanics, including shoulder flexion angles between $0^{\circ}$ and 2525$\circ$25 ∘ and abduction angles between $0^{\circ}$ and $20^{\circ}$ (Agha,, 2010; Altaboli et al.,, 2015; Gouvali and Boudolos,, 2006). Furthermore, keeping the arms on the desk significantly reduces strain on the spine. Consequently, according to ref (Poulakakis and Marmaras,, 1998)[50 Pheasant, 51 Poulakakis], the STH should be 3-5 cm higher than the SEH. A synchronized measure was created to verify that the SEH is the most important TH parameter and that the TH height should not increase above 50 mm above the SEH. Equation 5 expresses the link between TH and SEH.
$SEH\leq TH\leq(SEH+50)$ (5)
3.5.5 SSH against BH
Flexibility of the arms and upper trunk depends on the appropriate BH. It is recommended that the BH value is not higher than that of the scapula [31 Evans]. Therefore, an ideal BH is between 60% and 80% of sitting shoulder height (Agha,, 2010; Dianat et al.,, 2013; Castellucci et al.,, 2015). BH dimensions should also be less than or equal to the top margin of the scapula. This creates a match criterion, and Equation 6 shows how BH and SSH relate to one another.
$(0.60\times SSH)\leq BH\leq(0.80\times SSH)$ (6)
3.5.6 TT against SDC
The SDC should be sufficiently large to provide adequate space for leg movement and for the user to move the chair below the table (García-Acosta and Lange-Morales,, 2007; JFM et al.,, 2003). If the SDC is larger than the TT, it is considered a standard dimension (Parcells et al.,, 1999). However, the ideal dimension of SDC should be 20 mm higher than the knee height (Gouvali and Boudolos,, 2006). Therefore, a match criterion was developed, and the relationship between SDC and TT is shown in Equation 7.
$(TT+20)<SDC$ (7)
3.5.7 PH and TT against UTH
The user should have support for simple access to the seat, leg mobility, and getting off the seat from the Underneath Table Height (UTH) or Under Desktop Height (UDH). In addition, the design must incorporate SH, TT, and an additional 20 mm of space to allow for a shift in leg posture and the lowest standard dimension of UTH (Pérez-Gosende,, 2017). However, according to Gouvali and Boudolos, (2006), the highest dimension of UTH cannot be greater than the difference between the maximum dimension of TH and the table’s thickness, which is 30mm. Consequently, the link between UTH, PH, and TT yields a match equation, as shown in Equation 8.
$(SH+TT+30)\leq UTH\leq(SEH+[(PH+30)\cos(5^{\circ})]+0.1483AL-30)$ (8)
3.5.8 HB against BW
The HB is regarded as a key metric in the BW design process. The dimensions of BW have been considered by several authors (Thariq,, 2010). Taifa and Desai, (2017) has recently suggested that HB is considered a pertinent anthropometric measurement for BW design. Consequently, an equation was created, and Equation 9 illustrates the relationship between HB and BW:
$BW\geq HB$ (9)
3.5.9 BKL against TL
The TL is crucial for providing sufficient room for shifting lower body posture or movement. Previous research has suggested that TL should be higher than BKL (Pérez-Gosende,, 2017). Thus, Equation 10 defines the TL minimal limit.
$TL\geq BKL$ (10)
3.5.10 SEB against TD
The TD and SEB anthropometric measurements were connected. The partial widths of the AL and SEB should be included in the lower bound of the TD (Pérez-Gosende,, 2017). Sufficient room is needed for a table so that one can enter and leave the seat. TD thus permits elbow abduction at $20^{\circ}$ with a 20 mm margin (Afzan et al.,, 2012; Agha,, 2010; Altaboli et al.,, 2015; Gouvali and Boudolos,, 2006). Equation 11 considers these factors.
$(0.5SEB+0.342AL+20)\leq TD\leq EFL$ (11)
3.5.11 SCH against UEB
Another important anthropometric criterion for evaluating UEB is SCH (Bendix and Bloch,, 1986). Numerous studies have indicated that the scapula and arm are immobile if the SCH is smaller than the UEB (Castellucci et al.,, 2010; García-Acosta and Lange-Morales,, 2007; Gouvali and Boudolos,, 2006). Equation 12, therefore, depicts the link between SCH and UEB.
$UEB\leq SCH$ (12)
3.6 Level of Compatibility
Some equations have single-directional limitations, while others have dual constraints to assess how well participants’ anthropometric measurements match furniture dimensions. Match and mismatch categories were designed to evaluate compatibility in single-directional connections. Three categories were established for the dual-directional relationships: (a) match (where the anthropometric measurement falls between the limits), (b) low mismatch (where the anthropometric measurement falls below the maximum limit of the relationship), and (c) high mismatch (where the minimum limit of the relationship exceeds the anthropometric measurement).
4 Results and Discussion
4.1 Analysis of Anthropometric Measurements
Eleven anthropometric measures were used to compute descriptive statistics using Microsoft Excel 2021. The following values were estimated to create a standard database, as shown in Table 4: minimum (min), maximum (max), mean, standard deviation, and various percentile values (5th, 50th, and 95th percentiles). The anthropometric measurements of the male and female participants in this study were computed independently.
Table 4: Descriptive statistics of the anthropometric measurements taken for both male and female participants (in mm)
Measurements Gender Min Max Mean Standard Deviation Percentile
5th 50th 95th
PH Male 414 489 444.37 13.37 422.65 444.96 466.49
female 373 447 415.1 13.5 395.2 414.6 438.4
SEH Male 188 308 234.7 16.73 205.99 235.05 261.23
female 191 282 231.3 16.7 200.6 230.3 257.6
BPL Male 402 498 454.03 18.17 422.91 454.1 483.72
female 394 478 447.2 17.3 412.8 448.2 471
AL Male 331 394 364.1 10.77 346.35 363.86 382.29
female 322 366 342.4 9.6 327 342.1 355.9
TT Male 131 178 156.14 7.76 143.46 155.99 168.68
female 126 175 145.1 9.7 129.3 144.6 161.4
SSH Male 470 556 511.78 14.9 488.6 512.03 539.42
female 447 522 488 13.3 467 488.9 509.2
SCH Male 484 540 511.7 10.22 495.19 511.42 528.01
female 469 524 493.6 10.8 475.1 493.8 510.1
EFL Male 426 471 450.46 7.67 437.08 450.21 462.3
female 385 428 406.9 9.5 391.7 407.5 421.4
SEB Male 396 492 447.47 15.36 424.46 447.56 470.46
female 390 464 422.9 17.2 396.1 421.2 452.9
HB Male 328 382 350.53 9.4 335.31 350.23 365.34
female 344 390 366.2 8.5 353.3 367.2 379.4
BKL Male 496 548 521.47 9.54 505 522.23 536.1
female 475 546 509 13.7 489.2 509 529.9
The normal bell-shaped curves observed in Figures 4(a) and 4(b) indicate the reliability and consistency of the collected data. In statistics, a normal distribution is often considered a hallmark of reliable data, as it suggests that the measurements are evenly distributed around the mean, with fewer extreme values and more values clustered around the center. The presence of these curves in the data for both sexes indicates that the measurements were taken accurately and consistently without significant outliers or inconsistencies. This reliability is crucial in ergonomic research, as it ensures that the collected data accurately represent the population and can be used to make informed decisions about furniture design and ergonomics. Additionally, the normal distributions in both male and female datasets suggest that the collected measurements represent the general population rather than being skewed or biased toward one gender. This further enhances the reliability and validity of the dataset and supports its applicability in various settings and industries.
(a) Distribution of male anthropometric data
(b) Distribution of female anthropometric data
Figure 5: Anthropometric data showing normality (a bell-shaped curve) in the case of both (a) male and (b) female distributions.
The Spearman correlation method was used to analyze the correlation between all 11 anthropometric measurements. As depicted in Figure 6, the heatmap generated from the analysis represents positive correlations in red and negative correlations in blue. Notably, the elbow fingertip length exhibited a medium positive correlation with PH (0.44), AL (0.46), and SCH (0.43). Conversely, HB was moderately negatively correlated with EFL (-0.43). SEH and BPL showed negligible correlation with other anthropometric measurements, with values close to zero. The Spearman correlation heatmap values ranged from -1 to +1, indicating the strength and direction of the correlations between the variables.
Figure 6: Correlation between all 11 anthropometric measurements (for both male and female combined data).
4.2 Analysis of Existing Furniture Dimensions
Table 5 lists the 11 existing furniture dimensions of the computer furniture that were measured. Type-1 indicates a table with a non-adjustable chair, whereas type-2 indicates a table with an adjustable chair.
Table 5: Existing two types of furniture dimensions
Dimensions Type-1 Furniture set (mm) Type-2 Furniture set (mm)
SH 457.2 431.8 - 533.4
SW 393.7 457.2
SD 406.4 431.8
BH 304.8 304.8
BW 355.6 393.7
UEB 406.4 406.4
STH 241.3 228.6 - 330.2
STC 88.9 95.25 - 196.85
UTH 546.1 628.65
TL 482.6 457.2
TD 749.3 749.3
4.3 Mismatch Analysis
To evaluate the degree of mismatch between the current library furniture and body measurements, equations 2 through 12 were applied. A mismatch was found if the equation was not satisfied for equations with one-way relationships. Three classes were established to quantify the match or mismatch for equations with two-way relations. A match was found when the anthropometric measurement was in between the ranges. When the relation’s highest limit is less than the anthropometric measurement, a low mismatch is shown. A high mismatch is observed when the relation’s lowest limit is greater than the measurement. To find a mismatch, the writers looked into 11 aspects of computer furniture. Individual matches or mismatches were identified by comparing the dimensions of the furnishings with the body measurements of each participant. Finally, a percentage representing the measured match or mismatch for the sample population was given.
In Table 6, for the first type of furniture (non-adjustable chair with non-adjustable table), the mismatch level is notably higher for male body measurements than females in parameters, such as SH is 74.17%, SW is 59.92%, BW is 60.33% and SD is 5.75% higher whereas BH is 54.75%, and STH is 6.17% lower. Only one parameter, UEB, was found to be a 100% match for both genders. The results were particularly poor for the table-related parameters, indicating that furniture was poorly fitted for its users in these dimensions. All the parameters are compared in Figure 7 to better understand the scenario.
Table 6: Summary of the level of mismatch and match percentages of existing furniture dimensions of 1st type of furniture (non-adjustable chair and table)
Dimensions and anthropometry Gender Match (%) Lower mismatch (%) Upper mismatch (%) Total mismatch (%)
SH against PH Male 86.67 13.33 0 13.33
female 12.5 87.5 0 87.5
SW against HB Male 78.67 0 21.33 21.33
female 18.75 0 81.25 81.25
SD against BPL Male 92 8 0 8
female 86.25 13.75 0 13.75
BH against SSH Male 39 0 61 61
female 93.75 0 6.25 6.25
BW against HB Male 70.33 - - 29.67
female 10 - - 90
UEB against SCH Male 100 - - 0
female 100 - - 0
STH against SEH Male 66.33 0.33 33.33 33.67
female 72.5 1.25 26.25 27.5
STC against TT Male 0 -  - 100
female 0 -  - 100
UTH against TT, SEH, PH and AL Male 0 0 100 100
female 0 0 100 100
TL against BKL Male 0 - - 100
female 2.5 - - 97.5
TD against SEB, AL and EFL Male 0 100 0 100
female 0 100 0 100
Figure 7: Mismatch percentages of existing furniture dimensions of 1st type of furniture (non-adjustable chair and table).
The results for the second type of furniture (adjustable chair and non-adjustable table) in Table 7 show significantly lower mismatch levels than the first type of furniture for SH, SW, BW, STC, and UTH. For this particular furniture set, BW, UEB, and STH showed zero mismatches and needed no improvements. However, the outcomes for the table-related parameters remain subpar, indicating that furniture is still not well-suited for its users in these dimensions. Figure 8 compares all the parameters more explicitly.
Table 7: Summary of the level of mismatch and match percentages of existing furniture dimensions of 2nd type of furniture (adjustable chair and non-adjustable table)
Dimensions and anthropometry Gender Match (%) Lower mismatch (%) Upper mismatch (%) Total mismatch (%)
SH against PH Male 100 0 0 0
female 82.5 17.5 0 17.5
SW against HB Male 43 57 0 57
female 96.25 3.75 0 3.75
SD against BPL Male 49.67 50.33 0 50.33
female 38.75 61.25 0 61.25
BH against SSH Male 39 0 61 61
female 93.75 0 6.25 6.25
BW against HB Male 100 - - 0
female 100 - - 0
UEB against SCH Male 100 - - 0
female 100 - - 0
STH against SEH Male 100 0 0 0
female 100 0 0 0
STC against TT Male 99.33 - - 0.67
female 100 - - 0
UTH against TT, SEH, PH and AL Male 92 0 8 8
female 97.5 0 2.5 2.5
TL against BKL Male 0 - - 100
female 0 - - 100
TD against SEB, AL and EFL Male 0 100 0 100
female 0 100 0 100
Figure 8: Mismatch of existing furniture dimensions of 2nd type of furniture (adjustable chair and non-adjustable table).
4.4 ANOVA Test
A one-way ANOVA test was used to determine whether there were any statistically significant differences between the means of three or more independent groups. In this study, furniture dimensions were tested against anthropometric dimensions to determine if there were any significant differences. The level of significance ($\alpha$) value was set at 5%, meaning that a result was considered statistically significant if the p-value was $\leq 5\%$. The result was not statistically significant if the p-value was $>5\%$.
The ANOVA test resulted in a critical value (F-value) and a probable value (p-value). The F-value measures the variance between groups compared with the variance within the groups. The p-value is the probability of observing a result as extreme as that obtained, assuming that the null hypothesis is true. If the p-value was less than the $\alpha$ value (5% in this case), the null hypothesis was rejected, and it was concluded that there was a statistically significant difference between the groups. If the p-value was greater than the $\alpha$ value, the null hypothesis was not rejected, and it was concluded that there was no statistically significant difference between the groups.
In this study, ANOVA was used to determine whether there were any statistically significant differences in furniture dimensions and anthropometric measurements, both for males and females. The results showed that SH, SW, BH, UEB, STH, STC, and TL of furniture dimensions were rejected for non-adjustable chairs and tables, meaning there were statistically significant differences between the groups. In contrast, the SD and STH were accepted, indicating that there were statistically no significant differences between the groups (Table 8). For the adjustable chair non-adjustable table, we considered both the highest and lowest limits of SH, STH, and STC. If either the highest or lowest limit is accepted for males and females, we consider the parameter accepted. According to the ANOVA test for adjustable chairs with non-adjustable tables, only SH, SD, and STH dimensions were accepted. The rest of the dimensions, SW, BH, BW, UEB, STC, and TL were rejected (Table 9).
  • $H_{0}$(null hypothesis): There is no effect or difference between the anthropometric measurements and existing furniture dimensions.
  • $H_{1}$(alternative hypothesis): There is an effect or difference between the anthropometric measurements and existing furniture dimensions.
Table 8: ANOVA test results for non-adjustable furniture dimensions
Furniture Dimensions vs Anthropometric Dimensions Gender Value Type Percentile F -value p - value Decision
5th 50th 95th
SH vs PH Male Observed 422.65 444.96 466.49 0.71 0.447 Accept
Expected 443.79 457.2 470.61
Female Observed 395.17 414.58 438.4 7.83 0.049 Reject
Expected 443.79 457.2 470.61
SW vs HB Male Observed 335.31 350.23 365.34 11.19 0.029 Reject
Expected 376.98 393.7 410.42
Female Observed 353.28 367.24 379.45 4.51 0.101 Accept
Expected 375.97 393.7 411.43
SD vs BPL Male Observed 422.91 454.1 483.72 5.38 0.081 Accept
Expected 388.67 406.4 424.13
Female Observed 412.78 448.24 470.99 3.61 0.13 Accept
Expected 388.67 406.4 424.13
BH vs SSH Male Observed 488.6 512.03 539.42 173.83 0 Reject
Expected 294.62 304.8 314.98
Female Observed 467.05 488.85 509.16 184.73 0 Reject
Expected 294.62 304.8 314.98
BW vs HB Male Observed 335.31 350.23 365.34 0.28 0.625 Accept
Expected 346.88 355.6 364.32
Female Observed 353.28 367.24 379.45 1.48 0.29 Accept
Expected 346.88 355.6 364.32
UEB vs SCH Male Observed 495.19 511.42 528.01 70.93 0.001 Reject
Expected 392.32 406.4 420.48
Female Observed 475.12 493.8 510.12 44.56 0.003 Reject
Expected 392.32 406.4 420.48
STH vs SEH Male Observed 205.99 235.05 261.23 0.17 0.694 Accept
Expected 230.81 241.3 251.79
Female Observed 200.65 230.26 257.61 0.45 0.538 Accept
Expected 230.81 241.3 251.79
STC vs TT Male Observed 143.46 155.99 168.68 58.13 0.002 Reject
Expected 80.32 88.9 97.48
Female Observed 129.33 144.58 161.45 28.58 0.006 Reject
Expected 80.32 88.9 97.48
TL vs BKL Male Observed 505 522.23 536.1 13.8 0.021 Reject
Expected 473.68 482.6 491.52
Female Observed 489.25 509.05 529.86 4.37 0.105 Accept
Expected 473.68 482.6 491.52
Table 9: ANOVA test results for adjustable chair with non-adjustable table dimensions
Furniture Dimensions vs Anthropometric Dimensions Gender Value Type Percentile F-value p-value Decision
5th 50th 95th
SH (lowest limit) vs PH Male Observed 422.65 444.96 466.49 0.527 0.508 Accept
Expected 410.21 431.8 453.39
Female Observed 395.17 414.58 438.4 0.796 0.423 Accept
Expected 410.21 431.8 453.39
SH (highest limit) vs PH Male Observed 422.65 444.96 466.49 19.804 0.011 Reject
Expected 506.73 533.4 560.07
Female Observed 395.17 414.58 438.4 35.008 0.004 Reject
Expected 506.73 533.4 560.07
SW vs HB Male Observed 335.31 350.23 365.34 45.836 0.002  Reject
Expected 434.34 457.2 480.06
Female Observed 353.28 367.24 379.45 35.436 0.004 Reject
Expected 434.34 457.2 480.06
SD vs BPL Male Observed 422.91 454.1 483.72 1.023 0.369 Accept
Expected 410.21 431.8 453.39
Female Observed 412.78 448.24 470.99 0.337 0.593 Accept
Expected 410.21 431.8 453.39
BH vs SSH Male Observed 488.6 512.03 539.42 148.401 0 Reject
Expected 289.56 304.8 320.04
Female Observed 467.05 488.85 509.16 149.574 0 Reject
Expected 289.56 304.8 320.04
BW vs HB Male Observed 335.31 350.23 365.34 9.223 0.039 Reject
Expected 374.02 393.7 413.39
Female Observed 353.28 367.24 379.45 3.926 0.119 Accept
Expected 374.02 393.7 413.39
UEB vs SCH Male Observed 495.19 511.42 528.01 48.612 0.002 Reject
Expected 386.08 406.4 426.72
Female Observed 475.12 493.8 510.12 31.274 0.005 Reject
Expected 386.08 406.4 426.72
STH (lowest limit) vs SEH Male Observed 205.99 235.05 261.23 0.101 0.766 Accept
Expected 217.17 228.6 240.03
Female Observed 200.65 230.26 257.61 0.003 0.962 Accept
Expected 217.17 228.6 240.03
STH (highest limit) vs SEH Male Observed 205.99 235.05 261.23 26.745 0.007 Reject
Expected 313.69 330.2 346.71
Female Observed 200.65 230.26 257.61 28.057 0.006 Reject
Expected 313.69 330.2 346.71
STC (lowest limit) vs TT Male Observed 143.46 155.99 168.68 61.03 0.001 Reject
Expected 90.49 95.25 100.01
Female Observed 129.33 144.58 161.45 26.571 0.007 Reject
Expected 90.49 95.25 100.01
STC (highest limit) vs TT Male Observed 143.46 155.99 168.68 19.526 0.012  Reject
Expected 187.01 196.85 206.69
Female Observed 129.33 144.58 161.45 22.616 0.009 Reject
Expected 187.01 196.85 206.69
TL vs BKL Male Observed 505 522.23 536.1 16.011 0.016 Reject
Expected 434.34 457.2 480.06
Female Observed 489.25 509.05 529.86 8.739 0.042 Reject
Expected 434.34 457.2 480.06
4.5 Proposed Dimensions of Furniture
Autodesk Fusion 360 software was used for designing furniture. Table 10 outlines the suggested dimensions for furniture, categorized into chairs and tables, based on principles that consider the anthropometric measurements of the participants. The proposed UTH for both table types is 645mm as shown in Figure 9. The SH is suggested to be between 400 mm and 450 mm for adjustable chairs, based on the 5th percentile of female popliteal height with 30 mm shoe clearance (see Figure 9(a)). The recommended SH is 430 mm for non-adjustable chairs, as shown in Figure 9(b). The proposed SW is 425 mm, based on the 95th percentile of female HB. The recommended SD is 385 mm, determined by the 5th percentile of female BPL. BH is suggested to be 350 mm, considering the 5th percentile of female SSH. The recommended BW is 390 mm, based on the 95th percentile of male SEB. UEB is proposed to be 465 mm, considering the minimum subscapular height of female participants. The suggested UTH is 645 mm, considering the 5th percentile of male SEH.
Figure 9: Proposed UTH for both types of furniture in mm.
(a) Adjustable chair
(b) Non-adjustable chair
Figure 10: Proposed dimensions for both types of chairs (in mm): (a) adjustable chair, (b) non-adjustable chair.
Table 10: Proposed dimensions for both types of furniture sets
Dimensions For type-1 furniture set For type-2 furniture set
SH 430 400 - 450
SW 425 425
SD 385 385
BH 350 350
BW 390 390
UEB 465 465
STH 260 235 - 310
STC 200 95.25 - 200
UTH 645 645
For non-adjustable chairs and tables, Table 11 presents the mismatch percentages of the proposed furniture measurements. This is also shown graphically in Figure 11. The SH against PH showed 5% and 12.5% for males and females, respectively. The SW against HB showed zero mismatch for Male and only 1.25 lower mismatch for females. The SD against BPL showed 7.33% high mismatch and 0.33 % low mismatch for males, whereas 0% high mismatch and 1.25% low mismatch for females. It is necessary to reduce the high mismatch for females as much as possible while designing the chair’s SD. The BH, BW, UEB, and STC showed a 100% match for both males and females. It was not possible to design a single non-adjustable chair and table set for 100% match for both males and females. So, the dimensions were maintained to balance a reduced mismatch percentage for both males and females for STH and UTH dimensions. For adjustable chairs with tables, it was possible to match the 100% furniture dimensions for both males and females. SH and STH showed better results for both males and females due to the chair’s height adjustability. SW, SD, and UTH showed similar results to non-adjustable chairs with table furniture sets. Though adjustable chairs showed improved results in some scenarios, we considered both types of furniture to be financially feasible and avoided design complexity.
VDT should be placed according to recent studies; computer monitor placement should be based on the needs of the particular user and should be performed within a modest height range below eye level with a head tilt that is physiologically beneficial (Woo et al.,, 2016). OSHA (2008) suggests that monitors should be positioned 50–100 cm away from users while maintaining a viewing angle of $15^{\circ}$$20^{\circ}$ (Weidling and Jaschinski,, 2015).
Using a keyboard and mouse is a frequent task in a computer lab; it should be placed in the frequent or regular zone of the Barnes and Squires Work Envelope Model. This indicates that the keyboard and mouse should be placed within the 394 mm of table depth and 1194 mm of table length (Norman,, 2013).
Compared to the previous mismatch percentages in university furniture, the proposed dimensions exhibited reduced mismatches for both males and females. The suggested measurements are more compatible than the previous dimensions, leading to improved comfort levels and reduced risks of MSDs for users.
Table 11: Proposed furniture mismatch analysis for type 1 (non-adjustable chair and table)
Dimensions and anthropometry Gender Match (%) Low mismatch (%) High mismatch (%) Total mismatch (%)
SH against PH Male 95 0 5 5
female 87.5 12.5 0 12.5
SW against HB Male 100 0 0 0
female 98.75 1.25 0 1.25
SD against BPL Male 92.33 0.33 7.33 7.67
female 98.75 1.25 0 1.25
BH against SSH Male 100 0 0 0
female 100 0 0 0
BW against HB Male 100 - - 0
female 100 - - 0
UEB against SCH Male 100 - - 0
female 100 - - 0
STH against SEH Male 87.5 8.75 3.75 12.5
female 87.5 8.75 3.75 12.5
STC against TT Male 100 - - 0
female 100 - - 0
UTH against TT, SEH, PH and AL Male 100 0 0 0
female 98.75 0 1.25 1.25
Figure 11: Mismatch of existing furniture dimensions of 1st type of furniture (non-adjustable chair and non-adjustable table).
Table 12 presents the mismatch percentages of the proposed furniture measurements for adjustable chairs and tables. This is also shown in Figure 12. Compared to previous mismatch percentages in university furniture, the proposed dimensions exhibited reduced mismatches for both males and females. The suggested measurements are more compatible than the previous dimensions, leading to improved comfort levels and reduced risks of MSDs for users.
Table 12: Proposed furniture mismatch analysis for type-2 (adjustable chairs with non-adjustable tables) furniture set
Dimensions and anthropometry Gender Match (%) Low mismatch (%) High mismatch (%) Total mismatch (%)
SH against PH Male 100 0 0 0
female 100 0 0 0
SW against HB Male 100 0 0 0
female 98.75 0 1.25 1.25
SD against BPL Male 92.33 0.33 7.33 7.67
female 98.75 1.25 0 1.25
BH against SSH Male 100 0 0 0
female 100 0 0 0
BW against HB Male 100 - - 0
female 100 - - 0
UEB against SCH Male 100 - - 0
female 100 - - 0
STH against SEH Male 100 0 0 0
female 100 0 0 0
STC against TT Male 100 - - 0
female 100 - - 0
UTH against TT, SEH, PH and AL Male 100 0 0 0
female 98.75 0 1.25 1.25
Figure 12: Mismatch of existing furniture dimensions of 2nd type of furniture (adjustable chair and non-adjustable table).
5 Conclusions
This research comprehensively analyzes anthropometric measurements and their correlation with furniture dimensions in university settings. It proposes new dimensions for non-adjustable chairs and tables, as well as adjustable chairs and non-adjustable tables, to better suit the needs of students. The ANOVA test results demonstrated significant differences between the proposed and existing furniture dimensions, indicating the potential benefits of new measurements regarding comfort and ergonomics. The proposed dimensions showed reduced mismatch percentages for both males and females compared with existing furniture, suggesting a higher level of compatibility. This could improve comfort levels and lower the risk of MSDs among students who use the furniture. These findings highlight the importance of considering anthropometric measurements when designing furniture for educational environments.
The limitations of this research include the reliance on a specific population (university students), the exclusion of certain anthropometric measurements, and the lack of consideration of other factors influencing furniture design (e.g., cultural preferences and environmental constraints). Additionally, the proposed furniture dimensions may not be universally applicable and require further validation in other settings or populations. Future research could explore the implementation of the proposed dimensions in real-world classroom settings and assess their long-term impacts on student comfort and well-being. Additionally, further studies should investigate the effects of different furniture designs on student performance and productivity in academic settings.
This study provides a comprehensive analysis of the ergonomic dimensions of furniture, particularly chairs and tables, in an educational setting. This highlights the importance of designing furniture that caters to anthropometric measurements of university students, ensuring comfort and promoting a healthy posture. By collecting anthropometric data and correlating them with existing furniture dimensions, this study identifies mismatches and provides recommendations for optimized furniture design. The findings have significant practical implications, as they can guide the development of furniture that promotes better ergonomics and reduces the risk of MSDs among students. Moreover, this study’s methodology and insights can be applied beyond the educational setting, informing furniture design for various industries and environments. This underscores the importance of considering cultural preferences and environmental constraints when designing furniture. This study contributes to the growing body of ergonomics research and can potentially improve the well-being and productivity of individuals in diverse settings.
Appendix A NMQ Survey Questionnaire Format
Ergonomic Assessment
Serial no:
Age:
Gender:
Date of Assessment:
Question: Do you experience pain or discomfort in any of the following body parts when using your computer? If “yes,” please mention which part of your body.
(a) Neck (b) Shoulders/upper arm (c) Upper back (d) Lower back (e) Elbow/forearm (f) Wrist/Hand (g) Hips/Buttocks/Thighs (h) Knees and legs (i) Feet/Ankles
Please mention in which form is affected.
  1. (i)
    Constantly (most time of the day)
  2. (ii)
    Frequently (more than four times a month)
  3. (iii)
    Occasionally (two to four times a month)
  4. (iv)
    Never